# LEAP-VO: Long-term Effective Any Point Tracking for Visual Odometry

### v1

@@ 2 Related Work
In the realm of long-term point tracking or Tracking Any Point (TAP), various methods have been introduced recently. PIPs [14] reframes pixel tracking as a long-term motion estimation problem and significantly improves tracking accuracy by associating each pixel with a trajectory that spans multiple image frames. Context-TAP [2] and CoTracker [16] further tackle the spatial information by integrating spatial context features or tracking multiple queries together. TAP-Net [8] adopts a different approach that independently locates suitable candidate point correspondences with globally estimated cost volume decoded to correspondence position and occlusion, followed by TAPIR [9] which refines the global trajectory using a convolution layer based on TAP-Net’s predictions. Other approaches explore self-supervision targets to learn visual correspondence via powerful feature representation [15, 5]. Recently, OmniMotion [33] offers a pure test-time optimization approach that infers the point trajectory and occlusion directly from videos with optical flow predictions. However, its slow inference speed makes it unsuitable for real-time applications such as visual odometry.
@@
Visual Odometry (VO) can be categorized into indirect (feature-based) and direct methods. Indirect methods, such as MonoSLAM [7] and ORB-SLAM [23], mostly rely on feature extraction and matching to estimate the camera pose. Direct methods, conversely, skip feature extraction and matching and utilize intensity to optimize camera pose and 3D scene point positions based on photometric errors [10, 11, 34], showing better performance on low-texture regions where keypoint tracking can easily fail. However, they are more vulnerable to illumination changes and issues with rolling shutter cameras. Advancements in deep learning have spawned end-to-end monocular VO methods in supervised [35, 3, 38, 37, 31, 27] and unsupervised [44, 39, 25] settings. DeepVO [35] utilizes Recurrent Neural Networks to model sequential data, while SfMLearner [44] develops an unsupervised framework to learn the depth and camera motion from unlabeled monocular videos. TartanVO [37] proposes the first VO model that generalizes to multiple datasets and real-world scenarios. DROID-SLAM [30] follows learning-based optimization from RAFT [29] and proposes an end-to-end system, integrating flow, confidence, and geometric optimization via iterative GRU updates. The followed-up work DPVO [31] further cuts computational complexity by replacing dense feature tracking with sparse patch tracking.
Dynamic Track Detection.
In addition to enhancing accuracy, numerous studies have focused on the detection of dynamic objects within videos. Methods based on segmentation, such as DynaSLAM [1] and RCVD [18], employ semantic cues from pre-trained models to exclude dynamic regions. However, these segmentation-based methods are constrained to predefined object categories and struggle to capture real-world motion (for example, differentiating between a stationary and a moving car). Alternatively, DytanVO [27] and ParticleSfM [42] adopt a trajectory-based motion detection approach. This method addresses dynamic scenes using estimated point track information, thereby generating reliable camera trajectories in complex motion scenarios. Nonetheless, they depend on dense optical flow estimation to capture global motion patterns. Furthermore, dynamic tracks can be detected effectively using additional depth information from sensors or model estimations, as this allows for a 3D consistency check [6, 41].
@@ 3.1 Preliminary: Tracking Any Point
Tracking arbitrary pixels in an image sequence is traditionally akin to estimating optical flow, where each pixel is mapped to a corresponding location in a subsequent frame. However, two-view approaches often overlook the rich temporal context within the image sequence, leading to suboptimal outcomes. Recent advancements address this by revisiting the particle video concept and improving long-term point tracking through learning-based techniques [14, 9, 16]. These TAP methods can trace points over multiple frames and detect occlusions, significantly boosting tracking robustness under challenging scenarios.
Formulation. Given a sequence of $S$ consecutive RGB images $\mathbf{I}=[\mathbf{I}_{1},...,\mathbf{I}_{S}],\mathbf{I}_{s}\in\mathbb{R}^{3\times H\times W}$ as input, the goal of TAP is to trace a set of query points across these frames. For a specific query point with the 2D pixel coordinate $\mathbf{x}_{q}\in\mathbb{R}^{2}$ in frame $s_{q}$, TAP calculates its trajectory across all $S$ images as $\mathbf{X}=[\mathbf{x}_{1},...,\mathbf{x}_{S}],\mathbf{x}_{s}\in\mathbb{R}^{2}$, given $\mathbf{x}_{s_{q}}=\mathbf{x}_{q}$. Additionally, TAP estimates the visibility of each point throughout the sequence as $\mathbf{V}=[v_{1},...,v_{S}]$, where $v_{s}\in\{0,1\}$ is a binary label of point visibility indicating whether a point in frame $s$ is visible or occluded. The TAP formulation for a single query can be described as
@@ 3.3 Temporal Probability Modeling
Supervised learning applied to large-scale datasets has yielded notable advancements in learning-based point tracking methods. Nonetheless, trajectory estimation errors remain a concern, especially under challenging conditions like occlusions, motion blur, or textureless regions. Recognizing the varying reliability of trajectories, integrating an uncertainty estimation mechanism into the correspondence estimation process is advantageous. This is particularly crucial for quality-sensitive tasks such as visual odometry.
@@
Figure 4: LEAP-VO pipeline. Given a new image $\mathbf{I}_{t}$ received at time $t$, the feature extractor extracts new keypoints from $\mathbf{I}_{t}$. Then, all the keypoints from the latest $S_{LP}$-frame $\mathbf{I}_{t-S_{LP}+1:t}$ are tracked across all other frames within the current LEAP window, followed by a track filtering step to remove outliers. Finally, the local BA module is used on the current BA window to update the camera poses and 3D positions of the extracted keypoints. The colored arrows denote the moving direction of each module when a new image is received.
@@ 4.1 Datasets and Metrics
Replica. Replica [28] provides synthetic indoor environments for simulating the image-capturing process with a moving camera. We use the camera trajectories provided by [43] for the localization task, resulting in 16 camera trajectories derived from 8 static scenes. Each sequence comprises 900 frames with rendered RGB images. We use “Sequence 1” for the visual odometry evaluation.
@@
TartanAir Shibuya. The TartanAir Shibuya dataset [24] features dynamic scenes from two scenarios: “Standing Humans” and “Road Crossing”. The majority of the humans remain stationary in ”Standing Humans” sequences, while ”Road Crossing” sequences provide a more challenging setting where multiple humans are moving in different directions. Each sequence in the dataset includes 100 frames with more than 30 tracked moving humans.
@@
TartanVO [37] 0.290 0.092 1.303
DytanVO [27] 0.110 0.087 1.477
DPVO [31] 0.115 0.072 1.975
LEAP-VO (Ours) 0.089 0.066 1.250
@@ 4.2 Implementation details
We employ CoTracker [16], a recent PIPs variant, as our base TAP model. We train our model on the TAP-Vid-Kubric [8] training set, following [16]. We load the pre-trained CoTracker weights for the image feature extractor and train LEAP for 100,000 steps using 4 NVIDIA A100 GPUs in parallel. During training, we use $K=4$ steps for iterative refinements, $N=256$ for the number of queries, $S=8$ for the model window, and $S_{LP}=16$ for the tracking window. For inter-track anchor-based attention, the number of anchors is 64. The loss weights are set to $w_{1}=1.0,w_{2}=0.5,w_{3}=0.5$. In LEAP-VO, we perform 4 steps for each bundle adjustment with the Huber Loss for distance metric. The bundle adjustment window size $S_{BA}$ is set to 15. Track filtering parameters are set at $\gamma_{v}=0.9,\gamma_{d}=0.9,\gamma_{u}=0.8,\gamma_{track}=3$. Please refer to our supplementary materials for more details.
@@ 4.3 Visual Odometry Results
Replica. We compare our method with other baselines on the Replica dataset, which contains eight static scenes without any moving objects. Challenges arise in two test sequences that contain prolonged occlusions, leading to the potential camera tracking loss. As illustrated in Table 1, our approach outperforms other VO methods on all three metrics, showing effectiveness under static environments and the capability to handle occlusions. Further insights are available in Section 4.5. Notably, the classical methods, ORB-SLAM2 [22] and DynaSLAM [1], show high accuracy for camera pose tracking due to their re-localization capabilities during track losses.
@@
TartanAir-Shibuya. Following DynaVO [27], we further evaluate the VO performance using the dynamic TartanAir-Shibuya dataset. We assess the Average Trajectory Error (ATE) across seven diverse scenes, ranging from simple to challenging. As depicted in Table 3, our method consistently outperforms all other SLAM and VO baselines in the majority of these scenes. Moreover, trajectory visualizations in Figure 5 also highlight the robust camera tracking capabilities of our method in dynamic environments.
@@ 4.5 Ablation Study
Track Filtering. To highlight the efficacy of our track filtering module in LEAP-VO, we conduct an ablation study on different masks used in track filtering. We study three types of filtering criteria: visibility, dynamic track, and confidence filtering. Table 4 shows the impact of applying various filtering types on the dynamic MPI Sintel benchmark. Without dynamic filtering or confidence filtering, the performance of ATE drops by 45% and 70%, respectively. Combining all three types of filtering methods altogether, we achieve the best results in all three metrics.
@@
Anchor Sampling. Table 5 showcases the effectiveness of various anchor sampling strategies used in inter-track attention. Despite its distinctive features, SIFT sampling provides inferior performance due to its unpredictable keypoint distribution. In contrast, the random sampling tends to generate well-separated keypoints, leading to a performance similar to SIFT’s. Our distributed image gradient-based sampling shows the best performance, underscoring the significance of feature distribution and distinctiveness.
@@
Table 5: Comparison of anchor sampling policies ($N_{a}=64$) on TartanAir-Shibuya [24]. Our distributed image-gradient sampling method achieves better results compared to other policies.
@@ 5 Conclusion
Appendix A Implementation Details
A.1 LEAP
Training. We implement our approach in Python 3.10 with Pytorch 1.12. We train our model following the setting of CoTracker [16] on the TAP-Vid-Kubric [8] training set, which contains 11,000 synthetic sequences with 24 frames from MOVi-F dataset. Additionally, we extract the dynamic track label from the instance dynamic label provided by the Kubric simulator to supervise the dynamic track estimation. Notably, the MOVi-F dataset contains a limited number of scenes involving falling objects, accompanied by simple, linear camera motion at a constant velocity. In contrast, the TartanAir [36] training set, used by TartanVO [37], DytanVO [27], and DPVO [31], contains a wider range of environments and features more complex camera motions, along with a greater number of images. The image resolution of the synthetic sequences is $512\times 512$. During training, the input images are cropped to a resolution of $384\times 512$, and the feature map stride is 4. We adopt the AdamW [19] optimizer with the learning rate of $43e-$ 4 and linear warmup cosine annealing scheduler.
Model Architectures. For the image feature extractor, we employ the same Convolutional Neural Network (CNN) as used in PIPs [14]. It begins with a 7x7 convolution layer with a stride of 2, followed by an instance normalization layer and ReLU activation. The network then comprises several 3x3 kernel-sized residual blocks, each designed to process image features at different scales. The outputs of these layers are resized to a consistent scale using bilinear interpolation and then stacked. These stacked features are subsequently processed through 3x3 and 1x1 convolution layers. Regarding the refiner, we have explored integrating anchor-based inter-track attention with the MLP-based refiner from  [14] and the transformer-based refiner from  [16]. We find that both models achieve comparable point tracking accuracy given sufficient training time. However, the transformer-based network, which possesses five times more parameters, demonstrates faster convergence with fewer epochs of training. Consequently, we have chosen the transformer-based refiner for its training efficiency.
A.2 LEAP-VO
Keypoint Extration and Tracking. During the tracking process, we have introduced an additional hyperparameter, $S_{KF}$, to control the frequency of keypoint extraction, i.e., keypoints are extracted every $S_{KF}$ frame. The keypoints extracted within the most recent $S_{LP}$ are tracked bi-directionally within the LEAP window. A smaller $S_{KF}$ results in more frequent keypoint extraction, leading to a denser video representation and better capturing of motion patterns. However, extracting more keypoints also increases computational cost during the tracking front-end, especially due to the inter-track attention mechanism that exchanges information between every pair of tracks. We use $S_{KF}=2$ and $S_{LP}=12$ to balance performance and efficiency during the experiment.
Local Bundle Adjustment. In developing LEAP-VO, we build our system based on the local bundle adjustment (BA) implementation of DPVO [31]. The bundle adjustment window size, denoted as $S_{BA}$, is set to 15. This means that at each step, local BA optimizes the camera poses for the most recent 15 frames. Following local BA, we also incorporate the map outlier filtering process [26]. This process filters out points with large projection errors, specifically those where the distance between the projected correspondences and LEAP correspondences is substantial. Implementing this step helps to enhance the accuracy and robustness of the entire system during incremental tracking.
Initialization. We initiate the visual odometry system after collecting 12 frames, followed by 12 bundle adjustment (BA) steps to determine the initial camera pose and 3D scene points. As discussed in Section 3.4, keypoints are parameterized using their 2D-pixel coordinates from the extracted frame and a scalar depth value. Initially, we assign depth values by sampling from a uniform distribution within the range of $[0,1)$. After VO initialization, depth values are initialized using the median depth from the past three frames [31]. An additional advantage of our approach, which utilizes continuous feature tracking within a sliding window, is the ability to reuse estimated point locations from the previous window for better initialization in the current one. Let $\mathbf{X}_{i\rightarrow j}$ denote the estimated point position in frame $\mathbf{I}_{j}$, associated with its source frame $\mathbf{I}_{i}$. For all point tracking targets in the current local window $\mathbf{I}_{t-S_{LP}+1:t}$, it is observed that for all $i,j\in[t-S_{LP}+1,t-1]$, estimations have already been made by the previous local window $\mathbf{I}_{t-S_{LP}:t-1}$. Therefore, instead of initializing the point trajectory $\mathbf{X}^{0}$ by duplicating the query position $\mathbf{x}_{q}$, we can leverage the previous estimation for a better initialization.
Appendix B Additional Ablation Study
Temporal Probabilistic Distribution. In our Long-term Effective Any Point Tracking (LEAP) module, we employ a multivariate Cauchy distribution to model the temporal relationships among points from the same track. This approach is compared with the widely used multivariate Gaussian distribution. For this comparison, we train the LEAP network using the negative log-likelihood (NLL) loss with respect to the Gaussian distribution. The results of camera tracking using both LEAP (Cauchy) and LEAP (Gaussian) are presented in Table B.1. It is evident that LEAP-VO (Cauchy), with the LEAP (Cauchy) front-end, outperforms LEAP-VO (Gaussian) in all metrics. To further assess the differences, we evaluate the point tracking accuracy of both models on the TAP-Vid Davis First benchmark [8]. We use the standard metrics from [8], including the average Jaccard score (AJ), the average percentage of correct keypoints ($<\delta_{avg}^{x}$), and occlusion accuracy (OA), all calculated across various thresholds. The results, as shown in Table B.2, indicate that LEAP (Cauchy) surpasses LEAP (Gaussian) in point tracking accuracy, potentially leading to improved camera tracking performance in VO. We hypothesize that this improvement could be due to the L2 norm error term in the Gaussian NLL loss being less robust to outliers during training.
Method MPI Sintel
ATE RPE trans (m) RPE rot (deg)
LEAP-VO (Gaussian) 0.117 0.068 1.325
LEAP-VO (Cauchy) 0.089 0.066 1.250
Table B.1: Comparison of temporal distribution for camera tracking performance. Our Cauchy distribution formulation achieves better results compared to the Gaussian distribution.
Method TAP-Vid Davis (first)
AJ $<\delta^{x}_{avg}$ OA
LEAP (Gaussian) 50.855 66.032 83.177
LEAP (Cauchy) 56.889 73.301 85.005
Table B.2: Comparison of temporal distribution for point tracking performance. Our Cauchy distribution formulation achieves better results compared to the Gaussian distribution.
Appendix C More Visualization Results
Qualitative Results for Visual Odometry. We visualize the qualitative results for camera tracking, including dynamic track estimation, temporal probabilistic modeling, and comparisons of VO performance. Figure C.1 displays results from sample sequences of MPI-Sintel [4], TartanAir-Shibuya [24], and Replica [28]. In dynamic environments, methods that explicitly handle dynamic objects, such as DytanVO and ours, demonstrate superior performance compared to TartanVO, which does not address dynamic elements. DPVO, through implicitly learning the uncertainty from point correspondences, manages to handle dynamic scenes to some extent. However, it is not as effective as DytanVO and our method. This underscores the effectiveness of explicitly modeling dynamic elements in the formulation. Our method further surpasses DytanVO, showcasing the advantages of long-term point tracking over the two-view method.
Dynamic Track Estimation. We present additional visualizations of the dynamic track estimation results on the DAVIS dataset [17], as illustrated in Figure C.2. These visualizations demonstrate that our LEAP module is adept at handling dynamic scenes across a variety of scenarios.
Temporal Uncertainty Estimation. We present additional visualizations of the uncertainty estimation results on the MPI-Sintel [4] and Replica [28] datasets, as shown in Figure C.3 and Figure C.4, respectively. Guided by the NLL loss, our model tends to assign high uncertainty to areas with low texture and dynamic objects.
Figure C.1: Qualitative results for Visual Odometry on MPI-Sintel [4], TartanAir-Shibuya [24], and Replica [28]. Upper left: image sample with static (green) point tracking. Lower left: image sample with dynamic (red) and uncertain (yellow) point tracking. Right: comparison with the state-of-the-art VO methods.
Figure C.2: Visualization of dynamic track estimation on DAVIS [17]. From left to right: sample from image sequence, image with all point trajectories, all point trajectories, and estimated dynamic point trajectories.
Figure C.3: Visualization of point-wise uncertainty measurements on MPI-Sintel [4]. Keypoints with the lowest 20% (left) and highest 20% (right) uncertainty are shown in green and yellow, respectively.
Figure C.4: Visualization of point-wise uncertainty measurements on Replica [28]. Keypoints with the lowest 20% (left) and highest 20% (right) uncertainty are shown in green and yellow, respectively.
Appendix D Discussion and Future Work
Learning-based Anchor Selections. We have incorporated anchor-based methods into the LEAP pipeline to select points that capture motion patterns effectively. Our current method, which chooses anchors based on distributed image gradient sampling, outperforms uniform or random strategies. This technique selects easier-to-track anchors from high-gradient areas, ensuring comprehensive sequence coverage. However, it remains unclear whether this anchor selection strategy is optimal for the specific LEAP front-end. Exploring a learning-based anchor selection, integrated end-to-end with the point-tracking framework, could be beneficial. The main challenge lies in developing effective loss functions or self-supervised objectives for learning the anchor selection process.
Joint Refinement with Camera Tracking. In our visual odometry (VO) system, we distinctively handle the refinement of point trajectories and the refinement of camera pose and 3D map points as separate components. The refinement of point trajectories is managed using the refiner module in LEAP, which predicts iterative updates for point state variables including position, features, and other distribution parameters. For refining camera pose and 3D map points, we use a sliding-window local bundle adjustment module that leverages the Gauss-Newton method for updates. A promising future direction would be to merge point trajectory refinement with camera pose optimization into a unified refinement system.

### v2

@@ 2 Related Work
In the realm of long-term point tracking or Tracking Any Point (TAP), various methods have been introduced recently. PIPs [16] reframes pixel tracking as a long-term motion estimation with local information and improves tracking accuracy by associating each pixel with a trajectory that spans multiple image frames. The subsequent works enhance performance by integrating spatial context features [2], tracking multiple queries concurrently [18], and extending temporal fields and templates [49]. TAP-Net [9] adopts a different method by locating candidate correspondences through a global cost volume, improved by TAPIR’s trajectory refinement [10] and BootsTAP’s semi-supervised training [11]. Other approaches explore self-supervision targets to learn visual correspondence via powerful feature representation [17, 5, 36]. Recently, OmniMotion [37] provides a pure test-time optimization method that infers the point trajectory and occlusion directly from videos with optical flow predictions. However, its slow inference speed makes it unsuitable for real-time applications such as visual odometry.
@@
Visual Odometry (VO) can be categorized into indirect (feature-based) and direct methods. Indirect methods, such as MonoSLAM [8] and ORB-SLAM [25], mostly rely on feature extraction and matching to estimate the camera pose. Direct methods, conversely, skip feature extraction and matching and utilize intensity to optimize camera pose and 3D scene point positions based on photometric errors [12, 13, 38], showing better performance on low-texture regions where keypoint tracking can easily fail. However, they are more vulnerable to illumination changes and issues with rolling shutter cameras. Advancements in deep learning have spawned end-to-end monocular VO methods in supervised [39, 3, 30, 43, 40, 33, 28, 44] and unsupervised [51, 45, 27, 21] settings. DeepVO [39] utilizes Recurrent Neural Networks to model sequential data, while SfMLearner [51] develops an unsupervised framework to learn the depth and camera motion from unlabeled monocular videos. TartanVO [40] proposes a VO model that generalizes to multiple datasets and real-world scenarios. DROID-SLAM [32] follows learning-based optimization from RAFT [31] and proposes an end-to-end system, integrating flow, confidence, and geometric optimization via iterative GRU updates. The followed-up work DPVO [33] further cuts computational complexity by replacing dense feature tracking with sparse patch tracking.
Dynamic Track Estimation.
In addition to enhancing accuracy, numerous studies have focused on the detection of dynamic objects within videos. Methods based on segmentation, such as DynaSLAM [1] and RCVD [20], employ semantic cues from pre-trained models to exclude dynamic regions. However, these segmentation-based methods are constrained to predefined object categories and struggle to capture real-world motion (for example, differentiating between a stationary and a moving car). Alternatively, DytanVO [28] and ParticleSfM [48] adopt a trajectory-based motion detection approach. This method addresses dynamic scenes using estimated point track information, thereby generating reliable camera trajectories in complex motion scenarios. Nonetheless, they depend on dense optical flow estimation to capture global motion patterns. Furthermore, dynamic tracks can be detected using additional depth information for a 3D consistency check [7, 47], or by optimization based on photo consistency [35, 41, 42].
@@ 3.1 Preliminary: Tracking Any Point
Tracking arbitrary pixels in an image sequence is traditionally akin to estimating optical flow, where each pixel is mapped to a corresponding location in a subsequent frame. However, two-view approaches often overlook the rich temporal context within the image sequence, leading to suboptimal outcomes. Recent advancements address this by revisiting the particle video concept and improving long-term point tracking through learning-based techniques [16, 10, 18]. These TAP methods can track points over multiple frames and detect occlusions, significantly boosting tracking robustness under challenging scenarios.
Formulation. Given a sequence of $S$ consecutive RGB images $\mathbf{I}=[\mathbf{I}_{1},...,\mathbf{I}_{S}],\mathbf{I}_{s}\in\mathbb{R}^{3\times H\times W}$ as input, the goal of TAP is to track a set of query points across these frames. For a specific query point with the 2D pixel coordinate $\mathbf{x}_{q}\in\mathbb{R}^{2}$ in frame $s_{q}$, TAP predicts its trajectory across all $S$ images as $\mathbf{X}=[\mathbf{x}_{1},...,\mathbf{x}_{S}],\mathbf{x}_{s}\in\mathbb{R}^{2}$, given $\mathbf{x}_{s_{q}}=\mathbf{x}_{q}$. Additionally, TAP estimates the visibility of each point throughout the sequence as $\mathbf{V}=[v_{1},...,v_{S}]$, where $v_{s}\in\{0,1\}$ is a binary label of point visibility indicating whether a point in frame $s$ is visible or occluded. The TAP formulation for a single query can be described as
@@ 3.3 Temporal Probability Modeling
Supervised learning applied to large-scale datasets has yielded notable advancements in learning-based point tracking methods. Nonetheless, trajectory estimation errors remain a concern, especially under challenging conditions like occlusions, motion blur, or textureless regions. Recognizing the varying reliability of trajectories, integrating an uncertainty estimation mechanism into the correspondence estimation process is beneficial [6]. This is particularly crucial for quality-sensitive tasks such as visual odometry.
@@
Figure 4: LEAP-VO pipeline. Given a new image $\mathbf{I}_{t}$ received at time $t$, the keypoint extractor extracts new keypoints from $\mathbf{I}_{t}$. Then, all the keypoints from the latest $S_{LP}$-frame $\mathbf{I}_{t-S_{LP}+1:t}$ are tracked across all other frames within the current LEAP window, followed by a track filtering step to remove outliers. Finally, the local BA module is used on the current BA window to update the camera poses and 3D positions of the extracted keypoints. The colored arrows denote the moving direction of each module when a new image is received.
@@ 4.1 Datasets and Metrics
Replica. Replica [29] provides synthetic indoor environments for simulating the image-capturing process with a moving camera. We use the camera trajectories provided by [50] for the localization task, resulting in 16 camera trajectories derived from 8 static scenes. Each sequence comprises 900 frames with rendered RGB images. We use “Sequence 1" for the visual odometry evaluation.
@@
TartanAir Shibuya. The TartanAir Shibuya dataset [26] features dynamic scenes from two scenarios: “Standing Humans" and “Road Crossing". The majority of the humans remain stationary in "Standing Humans" sequences, while "Road Crossing" sequences provide a more challenging setting where multiple humans are moving in different directions. Each sequence in the dataset includes 100 frames with more than 30 tracked moving humans.
@@
TartanVO [40] 0.238 0.093 1.305
DytanVO [28] 0.131 0.097 1.538
DPVO [33] 0.076 0.078 1.722
LEAP-VO (Ours) 0.037 0.055 1.263
@@ 4.2 Implementation details
We employ CoTracker [18], a recent PIPs variant, as our base TAP model. We train our model on the TAP-Vid-Kubric [9] training set, following [18]. We load the pre-trained CoTracker weights for the image feature extractor and train LEAP for 100,000 steps using 4 NVIDIA A100 GPUs in parallel. During training, we use $K=4$ steps for iterative refinements, $N=256$ for the number of queries, $S=8$ for the model window, and $S_{LP}=12$ for the tracking window. For inter-track anchor-based attention, the number of anchors is 64. The loss weights are set to $w_{1}=1.0,w_{2}=0.5,w_{3}=0.5$. In LEAP-VO, we perform 4 steps for each bundle adjustment with the Huber Loss for distance metric. The bundle adjustment window size $S_{BA}$ is set to 15. Track filtering parameters are set at $\gamma_{v}=0.9,\gamma_{d}=0.9,\gamma_{u}=0.8,\gamma_{track}=3$. Please refer to our supplementary materials for more details.
@@ 4.3 Visual Odometry Results
Replica. We compare our method with other baselines on the Replica dataset, which contains eight static scenes without any moving objects. Challenges arise in two test sequences that contain prolonged occlusions, leading to the potential camera tracking loss. As illustrated in Table 1, our approach outperforms other VO methods on all three metrics, showing effectiveness under static environments and the capability to handle occlusions. Further insights are available in Section 4.5. For classical methods, ORB-SLAM2 [24] and DynaSLAM [1], we observe tracking loss and resuming via re-localization, which reduces long-term drift and shows high accuracy for camera pose tracking.
@@
TartanAir-Shibuya. Following DytanVO [28], we further evaluate the VO performance using the dynamic TartanAir-Shibuya dataset. We assess the Average Trajectory Error (ATE) across seven diverse scenes, ranging from simple to challenging. As depicted in Table 3, our method consistently outperforms all other SLAM and VO baselines in the majority of these scenes. Moreover, trajectory visualizations in Figure 5 also highlight the robust camera tracking capabilities of our method in dynamic environments.
@@ 4.5 Ablation Study
Track Filtering. To highlight the efficacy of our track filtering module in LEAP-VO, we conduct an ablation study on different masks used in track filtering. We study three types of filtering criteria: visibility, dynamic track, and confidence filtering. Table 4 shows the impact of applying various filtering types on the dynamic MPI Sintel benchmark. Without dynamic filtering or confidence filtering, the ATE increases by 46% and 130%, respectively. Combining all three types of filtering methods altogether, we achieve the best results in all three metrics.
@@
Keypoint Extraction. Table 5 showcases the effectiveness of various keypoint extraction strategies for VO performance. Despite its distinctive features, SIFT extraction provides inferior performance due to its unpredictable keypoint distribution. In contrast, the random sampling tends to generate well-separated keypoints, leading to a performance similar to SIFT’s. Our distributed image gradient-based sampling shows the best performance, underscoring the significance of feature distribution and distinctiveness.
@@
Table 5: Comparison of keypoint extraction policies ($N_{a}=64$) on TartanAir-Shibuya [26]. Our distributed image-gradient sampling method achieves better results compared to other policies.
Baseline comparison. We evaluate the pretrained CoTracker [18] for the end VO performance. Table 6 shows that LEAP front-end outperforms CoTracker on all datasets, demonstrating the effectiveness of the proposed dynamic track estimation and temporal probability modeling.
Front-end Replica MPI Sintel Shibuya
ATE RPE trans RPE rot ATE RPE trans RPE rot ATE
CoTracker 0.301 0.031 2.043 0.126 0.071 1.366 0.160
LEAP 0.204 0.030 1.992 0.037 0.055 1.263 0.029
Table 6: Baseline comparison with CoTracker front-end [18]. Our LEAP front-end achieves better results on all datasets.